OpenStudy (anonymous):

The product of some negative number and 4 less than twice that number = 336. find the number

OpenStudy (anonymous):

x(2x-4)=336

OpenStudy (a_clan):

x(2x-4) =336

OpenStudy (anonymous):

x^2-2x-168=0 (x-14)(x+12)=0 x=-12

OpenStudy (anonymous):

x^2-2x-168=0 (x-14)(x-12)=0 x=14 or -12 since we want a negative number, -12 -12 -28

OpenStudy (anonymous):

x(2x-4)=336 x(x-2)=168 x^2-2x-168=0

OpenStudy (anonymous):

ok - I was trying to do -x+2x-4=336 Thank you very much for your help everyone :-)